At what frequency will a 6.60-?F capacitor have a reactance of 2.30k? ?


Express your answer with the appropriate units.
Answer #1

Xc= 1/w*c ...w= 2*pie* frequency ...c= capacitance...so...


frequency=1/c*2*pie*X    =10.48 hz
Answer #2

2.3=1/2pif *c


c= 0.0953304 F


C=95.33 mF
Answer #3


capacitive reactance

here f is the frequency in hertz and C is the capacitance in farads. X is the reactance of the capacitor.So,


x = 2.30 x 10^3 ohms

C= 6.6 x 10^-6 F


So, 2.30 x 10^3 = 1/ (2 pi x 6.6 x 10^-6 x f)


or, f = 1/ (2 pi x 6.6 x 10^-6 x2.30 x 10^3 )


= 1/ (2 pi x 6.6 x 10^-6 x2.30 x 10^3 )

= 1/ 95.3304 x 10^-3


= 10.489 HZ
Answer #3


capacitive reactance

here f is the frequency in hertz and C is the capacitance in farads. X is the reactance of the capacitor.So,


x = 2.30 x 10^3 ohms

C= 6.6 x 10^-6 F


So, 2.30 x 10^3 = 1/ (2 pi x 6.6 x 10^-6 x f)


or, f = 1/ (2 pi x 6.6 x 10^-6 x2.30 x 10^3 )


= 1/ (2 pi x 6.6 x 10^-6 x2.30 x 10^3 )

= 1/ 95.3304 x 10^-3


= 10.489 HZ